Let X = {a, b, c} defined by f : X ®X such that f = {(a, b), (b, a), (c,c)}.  Find the values of

f–1, f2 and f4

1
Expert's answer
2021-09-14T06:09:26-0400

"X = {a,b,c} \\
f={(a,b), (b,a), (c,c)} \\
f(a) = b \\
f(b) = a \\
f(c)=c \\
f^{-1} = {(b,a), (a,b), (c,c)} \\
f^2 = ff(a) =f(b) =a \\
ff(b) = f(a) = b \\
ff(c) =f(c) =c \\
f^2 = {(a,a), (b,b), (c,c)} \\
f^4 = {(a,a), (b,b), (c,c)}"